OpenStudy (anonymous):

Help? Multiply 5SqrT(8)*4SqrT(7) A)16SqrT(10) B)20SqrT(15) C)40SqrT(14) D) 40

OpenStudy (espex):

What steps have you taken so far?

OpenStudy (anonymous):

ive tried multiplying the outers and i got 20 of course, But if i multiply the inside its 56.

OpenStudy (espex):

Okay, so you have 56 under the radical, what are all the factors for 56? Can you get a square or two out of it?

OpenStudy (anonymous):

7.5 is the rounded number i got for the sqrt.

OpenStudy (espex):

I like to lay it out like this: 56 = 1*2*2*2*7 and then look to see if I can make a square to pull out.

OpenStudy (espex):

Notice what you have here:\[\sqrt{56}\rightarrow\sqrt{1*2*2*2*7}\rightarrow\sqrt{1*2^{2}*2*7}\rightarrow 2\sqrt{1*2*7}\]

OpenStudy (anonymous):

okay.

OpenStudy (espex):

But does it make sense to you?

OpenStudy (anonymous):

it does until we get to the last part. how do you ger the 2 outside the randical

OpenStudy (espex):

Well you know that you can break the radical up, so couldn't I just as easily have written the middle step as: \[\sqrt{1}*\sqrt{2^{2}}*\sqrt{2}*\sqrt{7}\]

OpenStudy (anonymous):

okay.

OpenStudy (espex):

Does that make sense?

OpenStudy (anonymous):

Yes Sir

OpenStudy (espex):

Excellent. :)

OpenStudy (espex):

Looks like you have this figured out. :)

OpenStudy (anonymous):

But how do i get the answer?

OpenStudy (espex):

You already have the answer. All we did was simplify the radical portion of your answer. So what do you get when you put the pieces back together?